OpenStudy (anonymous):

How do i solve y/5+2=-8

OpenStudy (bahrom7893):

y + 10 = -40 y = -50

OpenStudy (anonymous):

y/5+2=-8 y + 10 = -40 ->y = -40 -10 = -50